OpenStudy (anonymous):

how do you go about doing this equation cos1° +cos2° +cos3° +...+cos358° +cos359° = ?

OpenStudy (anonymous):

HINT: \( \cos 1 ^\circ + \cos 179^\circ = 0 \) , Since \( \cos (180^\circ -1)= -\cos 1^circ \)

OpenStudy (anonymous):

\[ \cos 179^\circ = \cos (180^\circ -1^\circ)= -\cos 1^\circ\]

OpenStudy (anonymous):

So with that everything cancels out leaving only \(\cos 180^\circ\) which is \( -1\), i.e your final answer.

OpenStudy (anonymous):

Yeah that makes sense to me now thank you!

OpenStudy (anonymous):

Glad to help.